# ReviveDiff: A Universal Diffusion Model for Restoring Images in Adverse Weather Conditions

### v2

@@ Abstract
Images captured in challenging environments–such as nighttime, smoke, rainy weather, and underwater–often suffer from significant degradation, resulting in a substantial loss of visual quality. The effective restoration of these degraded images is critical for the subsequent vision tasks. While many existing approaches have successfully incorporated specific priors for individual tasks, these tailored solutions limit their applicability to other degradations. In this work, we propose a universal network architecture, dubbed “ReviveDiff”, which can address various degradations and bring images back to life by enhancing and restoring their quality. Our approach is inspired by the observation that, unlike degradation caused by movement or electronic issues, quality degradation under adverse conditions primarily stems from natural media (such as fog, water, and low luminance), which generally preserves the original structures of objects. To restore the quality of such images, we leveraged the latest advancements in diffusion models and developed ReviveDiff to restore image quality from both macro and micro levels across some key factors determining image quality, such as sharpness, distortion, noise level, dynamic range, and color accuracy. We rigorously evaluated ReviveDiff on seven benchmark datasets covering five types of degrading conditions: Rainy, Underwater, Low-light, Smoke, and Nighttime Hazy. Our experimental results demonstrate that ReviveDiff outperforms the state-of-the-art methods both quantitatively and visually.
@@ I Introduction
Over the years, substantial progress has been made in addressing various image degradation challenges, including denoising [1], dehazing [2], dedarkening [3], and deraining [4]. These methods typically rely on specific priors tailored to individual tasks. For instance, Retinex theory [5] has been successfully employed in many low-light enhancement techniques [3], while the atmospheric scattering model [6] has been commonly applied to guide the dehazing efforts. Fourier [7] and wavelet transformations have been extensively employed to address issues such as low resolution, blurring, and noise. However, these existing approaches are largely specialized, focusing on singular tasks with highly tailored solutions. This specialization, while effective within narrow scopes, limits their applicability across different types of degradation.
@@
As research has progressed, more complex challenges, such as nighttime image dehazing [6, 8, 9] and underwater image enhancement [10, 11, 12, 13], have come to the forefront. Although these solutions have shown impressive achievements by employing techniques like Atmospheric Point Spread Function-guided Glow Rendering and posterior distribution processes for underwater color restoration, they remain constrained by their focus on specific degradation scenarios. Their tailored design potentially limits their applicability to other degradations.
@@
The root causes of image degradation vary significantly; for example, blurring is typically caused by motion, while noise and low resolution often stem from limitations in imaging equipment. In contrast, quality degradation under adverse conditions is predominantly due to natural phenomena such as fog, water, and low light. This type of degradation generally maintains the structural integrity of objects in the image, leading to less distortion and information loss than other degradation types. These characteristics have inspired us to leverage general image attributes, such as edges, noise levels, and color distortions to guide the distortion process across a wide range of adverse weather conditions, rather than relying on priors specific to an individual task.
@@
    We introduced a unique prior-guided loss function that ensures optimal pixel restoration by leveraging edge information to refine object shapes and structures while utilizing histogram information to guide accurate color correction under adverse conditions.
@@ II-A Image Restoration
Image restoration is a broad field aimed at recovering high-quality images from degraded inputs. Numerous single-task approaches have been proposed to tackle specific problems, such as deblurring, denoising, deraining, and super-resolution. For instance, MDARNet [17] employs dual attention for image deraining, and PREnet [18] focuses on progressive rain removal. In the context of dark conditions, Li et al. [19] introduced LLFormer with Fourier Transform for low-light restoration. Underwater image restoration includes MLLE [20], which uses locally adaptive contrast, and UTrans [10], which integrates multi-scale feature fusion. For nighttime enhancement, Jin et al. [6] developed a gradient-adaptive convolution , and Yan et al. [9] used high-low frequency decomposition. Although these methods demonstrate strong performance, each primarily addresses a single type of degradation or environment, limiting their applicability across multiple adverse conditions.
@@ II-B Diffusion Models in Vision
The diffusion models [29, 30] have emerged as powerful tools in the field of image generation and restoration and have achieved significant improvement. These models work by degrading a signal through Gaussian noise and then restoring it through a reverse process. Fei et al. [31] introduced a unified image restoration model that utilizes a diffusion prior. In another significant contribution, Luo et al. proposed IR-SDE [15], which focuses on image restoration through mean-reverting stochastic differential equations.
@@ III Methodology
As illustrated in Fig. 2, our ReviveDiff is a U-shaped based latent diffusion model, with stacked Coarse-to-Fine Blocks (C2FBlocks) and Multi-Attentional Feature Complementation (MAFC) modules to tackle diverse real-world image restoration tasks. Compared with other diffusion models used for image enhancement, our ReviveDiff performs fusion at both low-resolution, macro-level latent space and high-resolution, micro-level latent space from the original input for the decoding process and under the guidance of fine granularity.
@@ III-A Mean-Reverting Stochastic Differential Processes
Forward SDE: The forward SDE process is defined as the transformation of an input high-quality image into the Gaussian noise. Mathematically, this process is expressed as
@@ III-B Coarse-to-Fine Learning
In contrast to the Fine Branch, the Coarse Branch focuses on global information and operates on a significantly large receptive field of $31\times 31$, as depicted in Fig. 4. However, existing approaches, such as the non-local mechanism, vision transformer blocks, and convolutional layers with large kernel sizes, tend to be computationally expensive, consuming significant resources and slowing down processing speeds. To address this issue, we utilize grouped-dilated convolutions to efficiently expand the receptive field size without substantially increasing the computation load.
@@
This coarse-to-fine learning process enables our model to capture both broad contextual information and focused local details, ensuring a balanced and thorough restoration of images across different types of degradations. As experimental results show, by fusing coarse and fine features in latent space, our ReviveDiff achieves superior performance in a range of challenging image enhancement tasks.
@@ III-C Multi-Attentional Feature Complementation
LPIPS $\downarrow$ / / / / 0.028 / /
Methods MPRNET [40] MAXIM [23] MDARNet [17] SFNet [21] AWRaCLe [50] FMRNET [51] DA-CLIP [26]
PSNR $\uparrow$ 36.40 38.06 35.68 38.21 35.71 37.81 30.10
SSIM $\uparrow$ 0.965 0.977 0.961 0.974 0.966 0.974 0.918
LPIPS $\downarrow$ 0.077 0.048 / / / / /
Methods DiffUIR [25] DiffPlugin [27] PromptGIP [24] DCPT-Restormer [52] DCPT-NAFNet [52] DCPT-PromptIR [52] ReviveDiff (Ours)
PSNR $\uparrow$ 31.80 25.40 22.80 36.68 35.68 37.32 39.09
SSIM $\uparrow$ 0.932 0.698 0.662 0.975 0.973 0.978 0.979
LPIPS $\downarrow$ / / / / / / 0.012
@@ IV-A Datasets
Underwater Image Enhancement: For the underwater image enhancement task, we used three datasets: the UIEB dataset [55], the C60 [55] dataset, and the U45 dataset [56]. The UIEB dataset consists of 890 pairs of low-quality and high-quality images, while the C60 and U45 datasets comprise 60 and 45 challenging images, respectively, without reference images. Consistent with previous works, we split the UIEB dataset into 800 pairs for training and 90 images for testing (referred to as “T90”). We tested our ReviveDiff on all these three datasets.
@@
Image Desmoking: For the image desmoking task, we utilized the SMOKE [80] dataset, which consists of 132 paired real-world images captured in natural scenes with fog generated by a fog machine. 120 pairs of these images are used for training and the remaining 12 pairs for testing.
@@ IV-C Evaluation Metrics
In our experiments, we adopt several widely used evaluation metrics to comprehensively assess our model’s performance across various restoration tasks. We employ the Peak Signal-to-Noise Ratio (PSNR) [61] and the Structural Similarity Index Measure (SSIM) [62] to quantify the fidelity of the restored images compared to the ground truth. Additionally, the Learned Perceptual Image Patch Similarity (LPIPS) [89] metric is used to evaluate perceptual similarity, reflecting human visual perception. For tasks such as underwater image enhancement, where reference images may not be available, we further utilize no-reference metrics like the Natural Image Quality Evaluator (NIQE) [90] and the Underwater Color Image Quality Evaluation (UCIQE) [91] to assess naturalness and color accuracy. Together, these metrics provide a robust framework for evaluating both pixel-level accuracy and perceptual quality across diverse adverse conditions.
@@ IV-D1 Image Deraining
Images captured in rainy conditions often suffer from the detrimental effects of rain streaks. To evaluate our method’s deraining capability, we utilized three widely recognized metrics: PSNR, SSIM, and LPIPS. We compared our method with several SOTAs from 2016 to 2025.
As demonstrated in Table I, our ReviveDiff model achieves the highest scores across all these metrics. When compared to the second-best method, SFNet [21], our PSNR surpassed it by 0.88, making ReviveDiff the only method to achieve a PSNR over 39. This indicates the superior quantitative performance of our ReviveDiff in deraining. Visually, as depicted in Fig. 5, our results show a significant reduction in rain presence and are most closely aligned with the reference image in terms of quality.
@@ IV-D2 Underwater Image Enhancement
IV-D3 Low-light Image Enhancement
IV-D4 Nighttime Image Dehazing
@@ IV-D5 Real-world Image Desmoking
IV-E Ablation Studies
IV-E1 Effectiveness of Coarse-to-Fine Learning

### v3

@@ Abstract
Images captured in challenging environments–such as nighttime, smoke, rainy weather, and underwater–often suffer from significant degradation, resulting in a substantial loss of visual quality. The effective restoration of these degraded images is critical for the subsequent vision tasks. While many existing approaches have successfully incorporated specific priors for individual tasks, these tailored solutions limit their applicability to other degradations. In this work, we propose a universal network architecture, dubbed “ReviveDiff”, which can address various degradations and restore images to their original quality by enhancing and restoring their details. Our approach is inspired by the observation that, unlike degradation caused by movement or electronic issues, quality degradation under adverse conditions primarily stems from natural media (such as fog, water, and low luminance), which generally preserves the original structures of objects. To restore the quality of such images, we leveraged the latest advancements in diffusion models and developed ReviveDiff to restore image quality from both macro and micro levels across some key factors determining image quality, such as sharpness, distortion, noise level, dynamic range, and color accuracy. We rigorously evaluated ReviveDiff on seven benchmark datasets covering five types of degrading conditions: Rainy, Underwater, Low-light, Smoke, and Nighttime Hazy. Our experimental results demonstrate that ReviveDiff outperforms the state-of-the-art methods both quantitatively and visually.
@@ I Introduction
Over the years, substantial progress has been made in addressing various image degradation challenges, including denoising [1], dedarkening [2], and deraining [3]. These methods typically rely on specific priors tailored to individual tasks. For instance, Retinex theory [4] has been successfully employed in many low-light enhancement techniques [2], while the atmospheric scattering model [5] has been commonly applied to guide the dehazing efforts. Fourier [6] and wavelet transformations have been extensively employed to address issues such as low resolution, blurring, and noise. However, these existing approaches are largely specialized, focusing on singular tasks with highly tailored solutions. This specialization, while effective within narrow scopes, limits its applicability across different types of degradation.
@@
As research has progressed, more complex challenges, such as nighttime image dehazing [5, 7] and underwater image enhancement [8, 9, 10, 11], have emerged. Although these solutions have achieved impressive results by employing techniques such as Atmospheric Point Spread Function-guided Glow Rendering and posterior distribution processes for underwater color restoration, they remain constrained by their focus on specific degradation scenarios. Their tailored design potentially limits their applicability to other degradations.
@@
The root causes of image degradation vary significantly; for example, blurring is typically caused by motion, while noise and low resolution often stem from limitations in imaging equipment. In contrast, quality degradation under adverse conditions is predominantly due to natural phenomena such as fog, water, and low light. This type of degradation generally maintains the structural integrity of objects in the image, leading to less distortion and information loss than other degradation types. These characteristics have inspired us to leverage general image attributes, such as edges, noise levels, and color distortions, to guide the distortion process across a wide range of adverse weather conditions, rather than relying on priors specific to an individual task.
@@
    We introduced a unique prior-guided loss function that ensures optimal pixel restoration by leveraging edge information to refine object shapes and structures, while utilizing histogram information to guide accurate color correction under adverse conditions.
@@ II-A Image Restoration
Image restoration is a broad field aimed at recovering high-quality images from degraded inputs. Numerous single-task approaches have been proposed to tackle specific problems, such as deblurring, denoising, deraining, and super-resolution. For instance, MDARNet [15] employs dual attention for image deraining, and PREnet [16] focuses on progressive rain removal. In the context of dark conditions, Li et al. [17] introduced LLFormer with the Fourier Transform for low-light restoration. Underwater image restoration includes MLLE [18], which uses locally adaptive contrast, and UTrans [8], which integrates multi-scale feature fusion. For nighttime enhancement, Jin et al. [5] developed a gradient-adaptive convolution , and Yan et al. [7] used high-low frequency decomposition. Although these methods demonstrate strong performance, each primarily addresses a single type of degradation or environment, limiting their applicability across multiple adverse conditions.
@@ II-B Diffusion Models in Vision
The diffusion models [27] have emerged as powerful tools in the field of image generation and restoration and have achieved significant improvement. These models work by degrading a signal through Gaussian noise and then restoring it through a reverse process. Fei et al. [28] introduced a unified image restoration model that utilizes a diffusion prior. In another significant contribution, Luo et al. proposed IR-SDE [13], which focuses on image restoration through mean-reverting stochastic differential equations.
@@ III Methodology
As illustrated in Fig. 2, our ReviveDiff is a U-shaped latent diffusion model, with stacked Coarse-to-Fine Blocks (C2FBlocks) and Multi-Attentional Feature Complementation (MAFC) modules to tackle diverse real-world image restoration tasks. Compared with other diffusion models used for image enhancement, our ReviveDiff performs fusion at both low-resolution, macro-level latent space and high-resolution, micro-level latent space from the original input for the decoding process, and under the guidance of fine granularity.
@@ III-A Mean-Reverting Stochastic Differential Processes
Forward SDE: The forward SDE process is defined as the transformation of an input high-quality image into Gaussian noise. Mathematically, this process is expressed as
@@ III-B Coarse-to-Fine Learning
In contrast to the Fine Branch, the Coarse Branch focuses on global information and operates on a significantly larger receptive field of $31\times 31$, as depicted in Fig. 4. However, existing approaches, such as the non-local mechanism, vision transformer blocks, and convolutional layers with large kernel sizes, tend to be computationally expensive, consuming significant resources and slowing down processing speeds. To address this issue, we utilize grouped-dilated convolutions to efficiently expand the receptive field size without substantially increasing the computation load.
@@
This coarse-to-fine learning process enables our model to capture both broad contextual information and focused local details, ensuring a balanced and thorough restoration of images across different types of degradations. As experimental results show, by fusing coarse and fine features in the latent space, our ReviveDiff achieves superior performance in a range of challenging image enhancement tasks.
@@ III-C Multi-Attentional Feature Complementation
LPIPS $\downarrow$ 0.028
Methods MPRNet [37] MAXIM [21] AirNet [47] IR-SDE [13] MDARNet [15] PromptIR [48] M2PN [49]
PSNR $\uparrow$ 36.40 38.06 34.90 38.30 35.68 37.04 38.36
SSIM $\uparrow$ 0.965 0.977 0.966 0.981 0.961 0.979 0.985
LPIPS $\downarrow$ 0.077 0.048 0.014
Methods NDR [50] GANet  [51] FrePrompter [52] AnyIR [53] VLU-Net [54] Perceive-IR [55] ReviveDiff (Ours)
PSNR $\uparrow$ 38.33 37.53 38.75 37.99 38.60 38.41 39.09
SSIM $\uparrow$ 0.984 0.973 0.984 0.982 0.984 0.984 0.979
LPIPS $\downarrow$ 0.012
@@ IV-A Datasets
Underwater Image Enhancement: For the underwater image enhancement task, we used three datasets: the UIEB dataset [58], the C60 [58] dataset, and the U45 dataset [59]. The UIEB dataset consists of 890 pairs of low-quality and high-quality images, while the C60 and U45 datasets comprise 60 and 45 challenging images, respectively, without reference images. Consistent with previous works, we split the UIEB dataset into 800 pairs for training and 90 images for testing (referred to as “T90”). We tested our ReviveDiff on all three datasets.
@@
Image Desmoking: For the image desmoking task, we utilized the SMOKE [83] dataset, which consists of 132 paired real-world images captured in natural scenes with fog generated by a fog machine. 120 pairs of these images are used for training, and the remaining 12 pairs for testing.
@@ IV-C Evaluation Metrics
In our experiments, we adopt several widely used evaluation metrics to assess our model’s performance across various restoration tasks comprehensively. We employ the Peak Signal-to-Noise Ratio (PSNR) [64] and the Structural Similarity Index Measure (SSIM) [65] to quantify the fidelity of the restored images compared to the ground truth. Additionally, the Learned Perceptual Image Patch Similarity (LPIPS) [92] metric is used to evaluate perceptual similarity, reflecting human visual perception. For tasks such as underwater image enhancement, where reference images may not be available, we further utilize no-reference metrics like the Natural Image Quality Evaluator (NIQE) [93] and the Underwater Color Image Quality Evaluation (UCIQE) [94] to assess naturalness and color accuracy. Together, these metrics provide a robust framework for evaluating both pixel-level accuracy and perceptual quality across diverse adverse conditions.
@@ IV-D1 Image Deraining
Images captured in rainy conditions often suffer from the detrimental effects of rain streaks. To evaluate the deraining capability of our method, we utilized three widely used metrics: PSNR, SSIM, and LPIPS. We compared our method with several SOTAs from 2016 to 2025.
As shown in Table I, our ReviveDiff model achieves a competitive performance across all these metrics. When compared to the second-best method, FrePrompter [52], our PSNR surpassed it by 0.34, making ReviveDiff the only method to achieve a PSNR over 39. This indicates the superior quantitative performance of our ReviveDiff in deraining. Visually, as depicted in Fig. 5, our results show a significant reduction in rain presence and are most closely aligned with the reference image in terms of quality.
@@ IV-D2 Underwater Image Enhancement
IV-D3 Low-light Image Enhancement
IV-D4 Nighttime Image Dehazing
@@ IV-D5 Real-world Image Desmoking
IV-E Ablation Studies
IV-E1 Effectiveness of Coarse-to-Fine Learning